OpenStudy (anonymous):

]Which of the following represents x squared minus twenty−five all over x squared minus five x in simplified form? x minus five all over x x plus five all over x x plus five over x 1 minus five over x

OpenStudy (saifoo.khan):

Aroubaaaaa!

OpenStudy (aroub):

=x^2-25/x^2-5x = (x-5)(x+5)/x(x-5) =x+5/x x plus five over x

OpenStudy (aroub):

saifooo!! :D

OpenStudy (saifoo.khan):

;]